OpenStudy (anonymous):

What Are the Factors of 32&40 and what is the common factor

OpenStudy (anonymous):

32=2^5, factor is 2 40=2^3*5, factors are 2 and 5 so the common factor is 2

OpenStudy (anonymous):

Thanks & What is The Factors for 25 & 35 an what are there common factors

OpenStudy (anonymous):

even easier 25=5^2, 5 35=5*7, 5 & 7 5 is common factor

OpenStudy (anonymous):

Thankss

OpenStudy (anonymous):

What Does ^ Stand For